# Rotating black holes in de Rham-Gabadadze-Tolley massive gravity: Analytic calculation procedure

### v2

@@ VI Gauge freedom and the Field equations
$\{1-4(|Q|^{2}-m^{2}|S|^{2})-|E^{K}|^{2}\}[(\Delta_{r}^{K}E^{K}_{,r})_{,r}+(\Delta_{\mu}^{K}E^{K}_{,\mu})_{,\mu}]=-2(E^{K})^{*}[\Delta_{r}^{K}(E^{K}_{,r})^{2}+\Delta_{\mu}^{K}(E^{K}_{,\mu})^{2}].$ (155)
@@
$(1-\tilde{E}^{K}(\tilde{E}^{K})^{*})[(\Delta^{K}_{u}\tilde{E}^{K}_{,u})_{,u}+(\Delta^{K}_{v}\tilde{E}^{K}_{,v})_{,v}]$ $=-2(\tilde{E}^{K})^{*}[\Delta^{K}_{u}(\tilde{E}^{K}_{,u})^{2}+\Delta^{K}_{v}(\tilde{E}^{K}_{,v})^{2}].$
@@
$(1+C_{u}u)^{2}+C_{v}^{2}v^{2}=\frac{C_{u}^{2}}{M^{2}-M_{0}^{2}}\left[\left(\frac{\sqrt{M-M_{0}}}{C_{u}}+r-M\right)^{2}+\frac{C_{v}^{2}}{C_{u}^{2}}(M^{2}-M_{0}^{2})\mu^{2}\right]$ (173)
@@
$\tilde{\Psi}^{K}$ $=-\frac{\Delta_{r}-a^{2}\Delta_{\mu}}{r^{2}+a^{2}\mu^{2}},$ (180)
@@
$q$ $=a\frac{(r^{2}+a^{2})\Delta^{K}_{\mu}-(1-\mu^{2})\Delta_{r}}{(r^{2}+a^{2})^{2}\Delta^{K}_{\mu}-a^{2}(1-\mu^{2})^{2}\Delta^{K}_{r}},$ (186)
and $A=-Q_{*}\frac{ra(1-\mu^{2})}{\rho^{2}},B=-Q_{*}\frac{\mu(r^{2}+a^{2})}{\rho^{2}}$ and $A^{\prime}=-S_{*}\frac{ra(1-\mu^{2})}{\rho^{2}},B^{\prime}=-S_{*}\frac{\mu(r^{2}+a^{2})}{\rho^{2}}$.
X the non-homogeneous solution
@@
$\Delta_{r}^{\Lambda}=\frac{\partial\Delta_{r}}{\partial(m^{2}\Lambda)},~{}\Delta_{\mu}^{\Lambda}=\frac{\partial\Delta_{\mu}}{\partial(m^{2}\Lambda)},~{}\tilde{Z}^{\Lambda}=\frac{\partial\tilde{Z}}{\partial(m^{2}\Lambda)}.$ (190)
@@
$\Psi^{\Lambda}[(\Delta_{r}^{\Lambda}\tilde{Z}^{\Lambda}_{,r})_{,r}+(\Delta_{\mu}^{\Lambda}\tilde{Z}^{\Lambda}_{,\mu})_{,\mu}]$ $=\frac{1}{2}(\Psi^{\Lambda})^{2}(\Delta_{r,rr}^{\Lambda}+\Delta_{\mu,\mu\mu}^{\Lambda})+\Delta^{\Lambda}_{r}(\tilde{Z}^{\Lambda}_{,r})^{2}+\Delta^{\Lambda}_{\mu}(\tilde{Z}^{\Lambda}_{,\mu})^{2},$ (192)
@@
If we define $\mu=-\cos\theta$, the solutions transform into the Boyer-Lindquist coordinate system, which is
$ds^{2}=-\frac{\rho^{2}\Delta_{r}\Delta_{\theta}}{\Sigma^{2}}dt^{2}+\frac{\Sigma^{2}}{\rho^{2}}\left(d\varphi-\frac{(r^{2}+a^{2})\Delta_{\theta}-\sin^{2}\theta\Delta_{r}}{\Sigma^{2}}adt\right)^{2}+\frac{\rho^{2}}{\Delta_{r}}dr^{2}+\frac{\rho^{2}\sin^{2}\theta}{\Delta_{\theta}}d\theta^{2},$ (204)
@@
XI Conclusion and discussion
In this paper, we analytical derived the rotating black hole solutions in dRGT model. The black hole famaliy in dRGT massive gravity would contain 5 parameters: the Mass $M$, the charge $Q_{*}$, the angular momentum per unit mass $a=\frac{J}{M}$, the cosmological constant term $\Lambda$ and the Stückelberg charge $S_{*}$. When consider $\Lambda\sim 1$, the cosmological constant term would be equal to $m^{2}$. However, the Stückelberg charge $S_{*}$ would still amend the charge $Q_{*}$.
@@ XI Conclusion and discussion
where $f=1-\frac{2M}{r}-\frac{m^{2}S_{*}^{2}}{r^{2}}-\frac{m^{2}}{3}r^{2}$. In Ref. [14], we obtain the static solution with $f=1-\frac{2M}{r}-\frac{S^{2}}{r^{\lambda}}-\Lambda r^{2}$. We found that in the special case $\lambda=2$, it can be returned to the solution of this article. For other $\lambda$ values, the solution has no corresponding rotating form. Thus, the black holes in dRGT model are divided into two categories: one kind of black hole solution has spherically symmetric form only; the other solution can be extended from spherical form to axisymmetric form.
The solution (204) can also be derived using the Newman-Janis algorithm, which is discussed in another paper. Additionally, we have not provided the expression for the reference metric $f_{ab}$. You can utilize the metric (204) to obtain $\bar{T}^{\mathcal{K}}_{ab}$. Meanwhile, we already have the expressions for $T^{\mathcal{K}}_{ab}$ from equations (77) - (82) . By setting these two equal and solving for them, you can derive the expressions for $f_{ab}$. These six equations will uniquely determine the six unknowns. However, the resulting expressions are quite complex, so we will not list them here. Readers who are interested are encouraged to find them on their own.
Besides, it is unclear whether this solution is the only possible rotational solution. The simplification of $T^{(\mathcal{K})}{}^{a}{}_{b}$ do not have an unique form. Therefore, other rotation solutions might be permitted. This will be the focus of future discussions.
Acknowledgement This work is partially supported by the National Natural Science Foundation of China (NSFC U2031112) and the Scientific Research Foundation of Hunan University of Arts and Sciences (23BSQD237). We also acknowledge the science research grants from the China Manned Space Project with NO. CMS-CSST-2021-A06.
Appendix A The derivation process 1
@@
$\left(\frac{\Delta_{r}}{\Psi}B_{,r}\right)_{,r}+\left(\frac{\Delta_{\mu}}{\Psi}B_{,\mu}\right)_{,\mu}$ $=\frac{\Delta_{r}}{\Psi^{2}}A_{,r}(\Phi_{,r}+2\text{Im}(HH^{*}_{,r}-m^{2}VV^{*}_{,r}))$
$\frac{\Delta_{\mu}}{\Psi^{2}}A_{,\mu}(\Phi_{,\mu}+2\text{Im}(HH^{*}_{,\mu}-m^{2}VV^{*}_{,\mu})),$ (213)
$\left(\frac{\Delta_{r}}{\Psi}A_{,r}\right)_{,r}+\left(\frac{\Delta_{\mu}}{\Psi}A_{,\mu}\right)_{,\mu}$ $=-\frac{\Delta_{r}}{\Psi^{2}}B_{,r}(\Phi_{,r}+2\text{Im}(HH^{*}_{,r}-m^{2}VV^{*}_{,r}))$
$-\frac{\Delta_{\mu}}{\Psi^{2}}B_{,\mu}(\Phi_{,\mu}+2\text{Im}(HH^{*}_{,\mu}-m^{2}VV^{*}_{,\mu})),$ (214)
@@
$\left(\frac{\Delta_{r}BB_{,r}}{\Psi}\right)_{,r}$ $+\left(\frac{\Delta_{\mu}BB_{,\mu}}{\Psi}\right)_{,\mu}=\frac{\Delta_{r}B_{,r}^{2}}{\Psi}+\frac{\Delta_{\mu}B_{,\mu}^{2}}{\Psi}+B\left(\left(\frac{\Delta_{r}B_{,r}}{\Psi}\right)_{,r}+\left(\frac{\Delta_{\mu}B_{,\mu}}{\Psi}\right)_{,\mu}\right)$
@@
$\left(\frac{\Delta_{r}AA_{,r}}{\Psi}\right)_{,r}$ $+\left(\frac{\Delta_{\mu}AA_{,\mu}}{\Psi}\right)_{,\mu}=\frac{\Delta_{r}A_{,r}^{2}}{\Psi}+\frac{\Delta_{\mu}A_{,\mu}^{2}}{\Psi}-\frac{\Delta_{r}}{\Psi^{2}}AB_{,r}(\Phi_{,r}+2\text{Im}(HH^{*}_{,r}-m^{2}VV^{*}_{,r}))$
@@
$2\Psi[$ $(\Delta_{r}\text{Im}HH^{*}_{,r})_{,r}+(\Delta_{\mu}\text{Im}HH^{*}_{,\mu})_{,\mu}]=-\Delta_{r}|H|^{2}_{,r}(\Phi_{,r}+2\text{Im}(HH^{*}_{,r}-m^{2}VV^{*}_{,r}))$
$-\Delta_{\mu}|H|^{2}_{,\mu}(\Phi_{,\mu}+2\text{Im}(HH^{*}_{,\mu}-m^{2}VV^{*}_{,\mu}))+2\Delta_{r}\Psi_{r}\text{Im}HH^{*}_{,r}+2\Delta_{\mu}\Psi_{\mu}\text{Im}HH^{*}_{,\mu}.$ (219)
@@
$2\Psi[$ $(\Delta_{r}\text{Im}VV^{*}_{,r})_{,r}+(\Delta_{\mu}\text{Im}VV^{*}_{,\mu})_{,\mu}]=-\Delta_{r}|V|^{2}_{,r}(\Phi_{,r}+2\text{Im}(HH^{*}_{,r}-m^{2}VV^{*}_{,r}))$
$-\Delta_{\mu}|V|^{2}_{,\mu}(\Phi_{,\mu}+2\text{Im}(HH^{*}_{,\mu}-m^{2}VV^{*}_{,\mu}))+2\Delta_{r}\Psi_{r}\text{Im}HH^{*}_{,r}+2\Delta_{\mu}\Psi_{\mu}\text{Im}HH^{*}_{,\mu}.$ (221)
Appendix B The derivation process 2
@@
$\Psi\{[\Delta_{r}(\Phi_{,r}+2\text{Im}(HH^{*}_{,r}-VV^{*}_{,r}))]_{,r}+[\Delta_{\mu}(\Phi_{,\mu}+2\text{Im}(HH^{*}_{,\mu}-VV^{*}_{,\mu}))]_{,\mu}\}$ $=$
$2\Delta_{r}\Psi_{,r}(\Phi_{,r}+2\text{Im}(HH^{*}_{,r}-VV^{*}_{,r}))+2\Delta_{\mu}\Psi_{,\mu}(\Phi_{,\mu}+2\text{Im}(HH^{*}_{,\mu}$ $-VV^{*}_{,\mu})).$ (222)
@@
$\Psi[$ $(\Delta_{r}\Phi_{,r})_{,r}+(\Delta_{\mu}\Phi_{,\mu})_{,\mu}]=$
$\Delta_{r}\{\Phi_{,r}(2\Psi+|H|^{2}-m^{2}|V|^{2})_{,r}+2(\Psi+|H|^{2}-m^{2}|V|^{2})_{,r}\text{Im}(HH^{*}_{,r}-VV^{*}_{,r})\}$
$+\Delta_{\mu}\{\Phi_{,\mu}(2\Psi+|H|^{2}-m^{2}|V|^{2})_{,\mu}+2(\Psi+|H|^{2}-m^{2}|V|^{2})_{,\mu}\text{Im}(HH^{*}_{,\mu}-VV^{*}_{,\mu})\}.$ (223)
Appendix C The derivation process 3
@@
$\frac{1}{2}(\Delta_{r,rr}$ $+\Delta_{\mu,\mu\mu})-\left(\Delta_{r}\frac{\Psi_{,r}}{\Psi}\right)_{,r}-\left(\Delta_{\mu}\frac{\Psi_{,\mu}}{\Psi}\right)_{,\mu}-\frac{\Psi^{2}}{\Delta_{\mu}}q_{,r}^{2}-\frac{\Psi^{2}}{\Delta_{r}}q_{,\mu}^{2}$
$=\frac{2}{\Psi}[\Delta_{r}(|H_{,r}|^{2}-m^{2}|V_{,r}|^{2})+\Delta_{\mu}(|H_{,\mu}|^{2}-m^{2}|V_{,\mu}|^{2})].$ (226)

### v3

@@ 6 Gauge freedom and the Field equations
$\{1-4(|Q|^{2}-m^{2}|S|^{2})-|E^{K}|^{2}\}[(\Delta_{r}^{K}E^{K}_{,r})_{,r}+(\Delta_{\mu}^{K}E^{K}_{,\mu})_{,\mu}]$
$=-2(E^{K})^{*}[\Delta_{r}^{K}(E^{K}_{,r})^{2}+\Delta_{\mu}^{K}(E^{K}_{,\mu})^{2}].$ (155)
@@
$(1-\tilde{E}^{K}(\tilde{E}^{K})^{*})[(\Delta^{K}_{u}\tilde{E}^{K}_{,u})_{,u}+(\Delta^{K}_{v}\tilde{E}^{K}_{,v})_{,v}]$ $=$
$-2(\tilde{E}^{K})^{*}[\Delta^{K}_{u}(\tilde{E}^{K}_{,u})^{2}+\Delta^{K}_{v}(\tilde{E}^{K}_{,v})^{2}]$ .
@@
$(1+C_{u}u)^{2}+C_{v}^{2}v^{2}$ $=\frac{C_{u}^{2}}{M^{2}-M_{0}^{2}}\bigg{[}\left(\frac{\sqrt{M-M_{0}}}{C_{u}}+r-M\right)^{2}$
$+\frac{C_{v}^{2}}{C_{u}^{2}}(M^{2}-M_{0}^{2})\mu^{2}\bigg{]}$ (173)
@@
$\tilde{\Psi}^{K}$ $=-\frac{\Delta^{K}_{r}-a^{2}\Delta^{K}_{\mu}}{r^{2}+a^{2}\mu^{2}},$ (180)
@@
$q$ $=a\frac{(r^{2}+a^{2})\Delta^{K}_{\mu}-(1-\mu^{2})\Delta^{K}_{r}}{(r^{2}+a^{2})^{2}\Delta^{K}_{\mu}-a^{2}(1-\mu^{2})^{2}\Delta^{K}_{r}},$ (186)
and $A=-Q_{*}\frac{ra(1-\mu^{2})}{\rho^{2}},B=-Q_{*}\frac{\mu(r^{2}+a^{2})}{\rho^{2}}$ and $A^{\prime}=-S_{*}\frac{ra(1-\mu^{2})}{\rho^{2}},B^{\prime}=-S_{*}\frac{\mu(r^{2}+a^{2})}{\rho^{2}}$. If one choosing $\mu=-\cos\theta$, the homogeneous solutions can be expressed in the Boyer-Lindquist coordinate system
$ds^{2}=$ $-\frac{\rho^{2}\Delta^{K}_{r}\Delta^{K}_{\theta}}{\Sigma^{2}}dt^{2}+\frac{\rho^{2}}{\Delta^{K}_{r}}dr^{2}+\frac{\rho^{2}\sin^{2}\theta}{\Delta^{K}_{\theta}}d\theta^{2}$
$+\frac{\Sigma^{2}}{\rho^{2}}\bigg{(}d\varphi-\frac{(r^{2}+a^{2})\Delta^{K}_{\theta}-\sin^{2}\theta\Delta^{K}_{r}}{\Sigma^{2}}adt\bigg{)}^{2},$ (187)
where
$\rho^{2}$ $=r^{2}+a^{2}\cos^{2}\theta,$ (188)
$\Sigma^{2}$ $=(r^{2}+a^{2})^{2}\Delta^{K}_{\theta}-a^{2}\sin^{4}\theta\Delta^{K}_{r},$ (189)
$\Delta^{K}_{r}$ $=r^{2}-2Mr+a^{2}+Q_{*}^{2}-m^{2}S_{*}^{2},$ (190)
$\Delta^{K}_{\theta}$ $=\sin^{2}\theta.$ (191)
10 the non-homogeneous solution
@@
$\Delta_{r}^{\Lambda}=\frac{\partial\Delta_{r}}{\partial(m^{2}\Lambda)},~\Delta_{\mu}^{\Lambda}=\frac{\partial\Delta_{\mu}}{\partial(m^{2}\Lambda)},~\tilde{Z}^{\Lambda}=\frac{\partial\tilde{Z}}{\partial(m^{2}\Lambda)}.$
@@
$\Psi^{\Lambda}[(\Delta_{r}^{\Lambda}\tilde{Z}^{\Lambda}_{,r})_{,r}+(\Delta_{\mu}^{\Lambda}\tilde{Z}^{\Lambda}_{,\mu})_{,\mu}]$ $=\frac{1}{2}(\Psi^{\Lambda})^{2}(\Delta_{r,rr}^{\Lambda}+\Delta_{\mu,\mu\mu}^{\Lambda})$
$+\Delta^{\Lambda}_{r}(\tilde{Z}^{\Lambda}_{,r})^{2}$ $+\Delta^{\Lambda}_{\mu}(\tilde{Z}^{\Lambda}_{,\mu})^{2},$ (196)
@@
Plugging expressions (205) - (207) into Eq. (99), one can finally obtain the metric. If we define $\mu=-\cos\theta$, the solutions transform into the Boyer-Lindquist coordinate system, which is
$ds^{2}=$ $-\frac{\rho^{2}\Delta_{r}\Delta_{\theta}}{\Sigma^{2}}dt^{2}+\frac{\rho^{2}}{\Delta_{r}}dr^{2}+\frac{\rho^{2}\sin^{2}\theta}{\Delta_{\theta}}d\theta^{2}$
$+\frac{\Sigma^{2}}{\rho^{2}}\left(d\varphi-\frac{(r^{2}+a^{2})\Delta_{\theta}-\sin^{2}\theta\Delta_{r}}{\Sigma^{2}}adt\right)^{2},$ (208)
@@
The reference metric $f_{ab}$ is given in appendix D.
11 Conclusion and discussion
In this paper, we analytical derived the rotating black hole solutions in dRGT model. The black hole family in dRGT massive gravity would contain 5 parameters: the mass $M$, the charge $Q_{*}$, the angular momentum per unit mass $a=\frac{J}{M}$, the cosmological constant term $\Lambda$ and the Stückelberg charge $S_{*}$. Compared to the original Kerr-Newman-Ads solution, the Stückelberg charge $S_{*}$ serves as a new characteristic, or ”hair,” of the black hole. While we treat the charges $S_{*}$ and $Q_{*}$ similarly, they have entirely different physical significances. The charge $S_{*}$, which is associated with the mass of the graviton, does not engage in electromagnetic interactions; it solely influences the curvature of spacetime. When a point charge $q$ moves through this spacetime, it experiences only the gravitational force from $S_{*}$ and not any electromagnetic force. Importantly, the charge $S_{*}$ would still exist even in the absence of the electric charge $Q_{*}$. By setting $m=0$, the solution (208) reproduces the Kerr-Newman solution. This occurs because the field equations (100) - (104) reduce to those found in Einstein-Maxwell theory when analyzed in that limit.
@@ 11 Conclusion and discussion
where $f=1-\frac{2M}{r}-\frac{m^{2}S_{*}^{2}}{r^{2}}-\frac{m^{2}}{3}r^{2}$. In Ref. Li2016-1 , we obtain the static solution with $f=1-\frac{2M}{r}-\frac{S^{2}}{r^{\lambda}}-\Lambda r^{2}$. We found that in the special case $\lambda=2$, it can be returned to the solution of this article. For other $\lambda$ values, the solution has no corresponding rotating form. Thus, the black holes in dRGT model are divided into two categories: one kind of black hole solution has spherically symmetric form only; the other solution can be extended from spherical form to axisymmetric form.
Although we can use the lemma to avoid calculating the Stückelberg field $\phi^{a}$, we still cannot express $I^{ab}=g^{\alpha\beta}\partial_{\alpha}\phi^{a}\partial_{\beta}\phi^{b}$ without $\phi^{a}$. If someone wants to determine whether the process $I^{ab}$ involves a singularity, the lemma must be applied in reverse. In the axisymmetric case, reversing the lemma would lead to another six differential equations. Addressing these equations would be challenging. This is the primary reason we utilize the lemma to bypass the calculation of the Stückelberg field $\phi^{a}$. As $a\rightarrow 0$, the metric (208) reduces to the metric in Ref. Li2016-1 , which obtains the Stückelberg field $\phi^{a}$ and a non-singular $I^{ab}$. Thus, the invariant $I^{ab}$ does not possess any singularities. However, in the case of $a\neq 0$, the question of whether $I^{ab}$ is singular remains an unresolved problem.
The computational process of this paper originates from Chandrasekhar Chandrasekhar:1983 , which is based on a general rotating axisymmetric metric. By utilizing symmetry, the field equations can be simplified into the Ernst equations. The process is akin to addressing spherically symmetric problems, although it involves complex mathematical techniques. Additionally, since it does not rely on excessive assumptions, this method is also applicable to modified gravity theories. Another commonly used method for obtaining axisymmetric solutions is the Newman-Janis algorithm. The solution presented in (208) can also be derived using this algorithm, which is further discussed in the paper Li2025-1 .
Acknowledgement This work is partially supported by the National Natural Science Foundation of China (NSFC U2031112) and the Scientific Research Foundation of Hunan University of Arts and Sciences (23BSQD237), and by Hunan Provincial Department of Education Key Laboratory of Information Detection and Intelligent Processing Technology, and by Applied Characteristic Subject of Hunan Province ”Electronic Science and Technology”. We also acknowledge the science research grants from the China Manned Space Project with NO. CMS-CSST-2021-A06.
Appendix A: The derivation process 1
@@
$\left(\frac{\Delta_{r}}{\Psi}B_{,r}\right)_{,r}+\left(\frac{\Delta_{\mu}}{\Psi}B_{,\mu}\right)_{,\mu}=\frac{\Delta_{r}}{\Psi^{2}}A_{,r}(\Phi_{,r}+2\text{Im}(HH^{*}_{,r}$
$-m^{2}VV^{*}_{,r}))+\frac{\Delta_{\mu}}{\Psi^{2}}A_{,\mu}(\Phi_{,\mu}+2\text{Im}(HH^{*}_{,\mu}-m^{2}VV^{*}_{,\mu})),$ (217)
$\left(\frac{\Delta_{r}}{\Psi}A_{,r}\right)_{,r}+\left(\frac{\Delta_{\mu}}{\Psi}A_{,\mu}\right)_{,\mu}=-\frac{\Delta_{r}}{\Psi^{2}}B_{,r}(\Phi_{,r}+2\text{Im}(HH^{*}_{,r}$
$-m^{2}VV^{*}_{,r}))-\frac{\Delta_{\mu}}{\Psi^{2}}B_{,\mu}(\Phi_{,\mu}+2\text{Im}(HH^{*}_{,\mu}-m^{2}VV^{*}_{,\mu})),$ (218)
@@
$\left(\frac{\Delta_{r}BB_{,r}}{\Psi}\right)_{,r}+\left(\frac{\Delta_{\mu}BB_{,\mu}}{\Psi}\right)_{,\mu}=\frac{\Delta_{r}B_{,r}^{2}}{\Psi}+\frac{\Delta_{\mu}B_{,\mu}^{2}}{\Psi}$
$+B\left(\left(\frac{\Delta_{r}B_{,r}}{\Psi}\right)_{,r}+\left(\frac{\Delta_{\mu}B_{,\mu}}{\Psi}\right)_{,\mu}\right)$
@@
$\left(\frac{\Delta_{r}AA_{,r}}{\Psi}\right)_{,r}+\left(\frac{\Delta_{\mu}AA_{,\mu}}{\Psi}\right)_{,\mu}=\frac{\Delta_{r}A_{,r}^{2}}{\Psi}+\frac{\Delta_{\mu}A_{,\mu}^{2}}{\Psi}$
$-\frac{\Delta_{r}}{\Psi^{2}}AB_{,r}(\Phi_{,r}+2\text{Im}(HH^{*}_{,r}-m^{2}VV^{*}_{,r}))$
@@
$2\Psi[(\Delta_{r}\text{Im}HH^{*}_{,r})_{,r}+(\Delta_{\mu}\text{Im}HH^{*}_{,\mu})_{,\mu}]=-\Delta_{r}|H|^{2}_{,r}(\Phi_{,r}$
$+2\text{Im}(HH^{*}_{,r}-m^{2}VV^{*}_{,r}))-\Delta_{\mu}|H|^{2}_{,\mu}(\Phi_{,\mu}+2\text{Im}(HH^{*}_{,\mu}$
$-m^{2}VV^{*}_{,\mu}))+2\Delta_{r}\Psi_{r}\text{Im}HH^{*}_{,r}+2\Delta_{\mu}\Psi_{\mu}\text{Im}HH^{*}_{,\mu}.$ (223)
@@
$2\Psi[(\Delta_{r}\text{Im}VV^{*}_{,r})_{,r}+(\Delta_{\mu}\text{Im}VV^{*}_{,\mu})_{,\mu}]=-\Delta_{r}|V|^{2}_{,r}(\Phi_{,r}$
$+2\text{Im}(HH^{*}_{,r}-m^{2}VV^{*}_{,r}))-\Delta_{\mu}|V|^{2}_{,\mu}(\Phi_{,\mu}+2\text{Im}(HH^{*}_{,\mu}$
$-m^{2}VV^{*}_{,\mu}))+2\Delta_{r}\Psi_{r}\text{Im}HH^{*}_{,r}+2\Delta_{\mu}\Psi_{\mu}\text{Im}HH^{*}_{,\mu}.$ (225)
Appendix B: The derivation process 2
@@
$\Psi\{[\Delta_{r}(\Phi_{,r}+2\text{Im}(HH^{*}_{,r}-VV^{*}_{,r}))]_{,r}+[\Delta_{\mu}(\Phi_{,\mu}+2\text{Im}(HH^{*}_{,\mu}$
$-VV^{*}_{,\mu}))]_{,\mu}\}=2\Delta_{r}\Psi_{,r}(\Phi_{,r}+2\text{Im}(HH^{*}_{,r}-VV^{*}_{,r}))$
$+2\Delta_{\mu}\Psi_{,\mu}(\Phi_{,\mu}+2\text{Im}(HH^{*}_{,\mu}-VV^{*}_{,\mu})).$ (226)
@@
$\Psi[(\Delta_{r}\Phi_{,r})_{,r}+(\Delta_{\mu}\Phi_{,\mu})_{,\mu}]=\Delta_{r}\{\Phi_{,r}(2\Psi+|H|^{2}-m^{2}|V|^{2})_{,r}$
$+2(\Psi+|H|^{2}-m^{2}|V|^{2})_{,r}\text{Im}(HH^{*}_{,r}-VV^{*}_{,r})\}$
$+\Delta_{\mu}\{\Phi_{,\mu}(2\Psi+|H|^{2}-m^{2}|V|^{2})_{,\mu}$
$+2(\Psi+|H|^{2}-m^{2}|V|^{2})_{,\mu}\text{Im}(HH^{*}_{,\mu}-VV^{*}_{,\mu})\}.$ (227)
Appendix C: The derivation process 3
@@
$\frac{1}{2}(\Delta_{r,rr}+\Delta_{\mu,\mu\mu})-\left(\Delta_{r}\frac{\Psi_{,r}}{\Psi}\right)_{,r}-\left(\Delta_{\mu}\frac{\Psi_{,\mu}}{\Psi}\right)_{,\mu}-\frac{\Psi^{2}}{\Delta_{\mu}}q_{,r}^{2}$
$-\frac{\Psi^{2}}{\Delta_{r}}q_{,\mu}^{2}=\frac{2}{\Psi}[\Delta_{r}(|H_{,r}|^{2}-m^{2}|V_{,r}|^{2})+\Delta_{\mu}(|H_{,\mu}|^{2}-m^{2}|V_{,\mu}|^{2})].$ (230)